# Generalized cryptographic semi-regular sequences: A variant of Fröberg conjecture and a simple complexity estimation for Gröbner basis computation

### v1

@@ 1 Introduction
  • The sequence $(f_{1}^{h},\ldots,f_{m}^{h})\in(R^{\prime})^{m+1}$ is generalized cryptographic semi-regular.
Then, by the first assumption, it follows from Theorem 1 (1) that the inequality (1.0.2) holds for $I=\langle F^{h}\rangle_{R^{\prime}}$ and $\ell=y$, where we replace $R$ by $R^{\prime}$ and set $\sigma$ to be the identity. Moreover, under the second and third assumptions, we also have the complexity bound
$O\left(m\binom{n+D^{(n+1,m)}}{D^{(n+1,m)}}^{\omega}\right)$ (1.0.6)
by Theorem 1 (2). Note that this upper-bound is heuristically found in [21].
@@ 2 Preliminaries
Let $K$ be a field, and let $\overline{K}$ denote its algebraic closure. Let $n$ be a natural number, and let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring of $n$ variables over $K$. For $d\geq 0$, we denote by $R_{d}$ the homogeneous component of $R$ of degree $d$. The Krull dimension of a commutative ring $A$ is denoted by $\mathrm{Krull.dim}(A)$. For ideals $I$ and $J$ of $R$, their ideal quotient is defined as $(I:J)=\{f\in R\mid fJ\subset I\}$. The saturation of $I$ with respect to $J$ is $(I:J^{\infty}):=\{f\in R\mid fJ^{s}\subset I\mbox{ for some $s\geq 1$}\}$. In particular, the saturation of a homogeneous ideal $I$ with respect to the maximal homogeneous ideal $\mathfrak{m}=\langle x_{1},\ldots,x_{n}\rangle_{R}$ is denoted by $I^{\rm sat}=(I:\mathfrak{m}^{\infty})$, which is also homogeneous.
@@ 2.1 Hilbert regularity
Let $I$ be a homogeneous ideal of $R$. Let ${\rm HP}_{R/I}\in\mathbb{Q}[z]$ denote the Hilbert polynomial of $R/I$, namely there exists a non-negative integer $d_{0}$ such that ${\rm HF}_{R/I}(d)={\rm HP}_{R/I}(d)$ for any $d$ with $d\geq d_{0}$, where $\mathrm{HF}$ is the Hilbert function.
Definition 2.1.1 (Hilbert regularity (or index of regularity [13, Chapter IX, Section 3])).
@@
    There exists a non-negative integer $d_{0}$ such that ${\rm HF}_{R/I}(d)={\rm HF}_{R/I}(d_{0})$ (resp. ${\rm HF}_{R/I}(d)=0$) for any $d$ with $d\geq d_{0}$.
@@
The middle and the right hand sides are called the generalized degree of regularity of $I$ for $r\leq 1$, denoted by $\widetilde{d}_{\rm reg}(I)$ (cf. [29, Sect. 2.3]): When $r=0$, these are equal to the degree of regularity $d_{\rm reg}(I):=\min\{d:R_{d}=I_{d}\}$. Namely, we have
@@
We also note that ${\rm HS}_{R/I}$ is written as
${\rm HS}_{R/I}(z)=\dfrac{h_{R/I}(z)}{(1-z)^{r}}$
for some polynomial $h_{R/I}(z)\in\mathbb{Z}[z]$, which is called the $h$-polynomial of $R/I$. Then, we have ${\rm hilb}(I)=\deg h_{R/I}-r+1$, see [6, Prop. 4.1.12].
@@
where ‘${\rm reg}$’ means Castelnuovo-Mumford regularity, see [15, §20.5] for the definition. The equality holds if $R/I$ is Artinian, i.e., the Krull dimension of $R/I$ is zero.
Definition 2.1.4 (cf. [24, Section 3]).
@@
Lemma 2.1.5 ([5, Lemma 4.1]; see also [32, Proposition 5.4]).
Let $I$ be a homogeneous ideal of $R$, and let $\prec$ be the reverse lexicographic order on $R$ with $x_{n}\prec\cdots\prec x_{1}$. Then, the following are equivalent:
@@
    $\mathrm{Krull.dim}(R/\langle I,x_{n-r+1},\ldots,x_{n}\rangle)=0$.
@@
    $\mathrm{Krull.dim}(R/\langle\langle\mathrm{LM}_{\prec}(I)\rangle,x_{n-r+1},\ldots,x_{n}\rangle)=0$.
In the proof of [5, Lemma 4.1], the authors first proved the equivalence of the conditions (2), (3), and (4), and the implication (1) $\Rightarrow$ (2). The non-trivial part is that (2), (3), and (4) imply (1), which follows from [32, Proposition 5.4].
@@
$\langle I,x_{n-i+1},\ldots,x_{n}\rangle^{\rm sat}=(\langle I,x_{n-i+1},\ldots,x_{n}\rangle:x_{n-i}^{\infty})$
@@
For equivalent definitions, see [23, Section 2] for details. Note that $I$ is in SNP if and only if $\langle\mathrm{LM}_{\prec}(I)\rangle$ is in SNP. Hashemi also proved the following inequality:
Theorem 2.1.7 ([23, Proposition 4.7 and Theorem 4.17]).
@@
for the graded lexicographic ordering $\prec$ on $R$ with $x_{n}\prec\cdots\prec x_{1}$.
Lemma 2.1.8 ([24, Lemma 23] and [25, Proposition 3.2 (v)]).
@@
    $I$ is in quasi stable position with respect to the ordered variables $x_{1},\ldots,x_{n}$ (see e.g., [24, Definition 6] for definition).
@@
The equivalence of (1) and (2) is proved in [24, Lemma 23], and that of (2) and (3) follows from [25, Proposition 3.2 (v)]. ∎
@@
Theorem 2.1.9 ([24, Proof of Theorem 30]).
Assume that $r=1$. If $R/\langle I,x_{n}\rangle$ is Artinian (or one of the equivalent conditions given in Lemma 2.1.8 holds), then we have the inequality (1.0.1) for the graded reverse lexicographic ordering $\prec$ on $R$ with $x_{n}\prec\cdots\prec x_{1}$.
@@ 2.2 Generalized semi-regular sequences
Definition 2.2.2 ([1, Definition 3]; see also [14, Definition 1]).
@@
Definition 2.2.3 ([1, Definition 5], [3, Definition 5]; see also [14, Section 2]).
@@
Proposition 2.2.4 ([14, Proposition 1 (d)]; see also [3, Proposition 6]).
With the same notation as in Theorem 2.2.1, a sequence $(f_{1},\ldots,f_{m})$ of homogeneous polynomials in $R\smallsetminus K$ is cryptograpic semi-regular if and only if
@@
Here, we recall the following lemma:
Lemma 2.2.5 ([20, b) on page 121]).
@@
in the lexicographic sense, i.e., the first coefficient that differs, if any, is greater on the left hand side, see [20, page 118].
@@
To deal with the case where $R/I$ has Krull dimension one, we extend the notion of cryptgraphic semi-regular as follows:
Definition 2.2.7 ([29, Definition 2.3.3]).
@@
By our assumption that $\bm{F}$ is generalized cryptographic semi-regular, the equality (2.2.1) holds for $d=\widetilde{d}_{\rm reg}(I)$. For any $d^{\prime}$ with $d^{\prime}<\widetilde{d}_{\rm reg}(I)$, we have ${\rm HF}_{R/I}(d^{\prime})>0$, whence the $z^{d^{\prime}}$-coefficient of $\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}$ is also positive. Thus, the inequality (2.2.3) holds. ∎
@@
$\bm{F}$ is regular, equivalently
${\rm HS}_{R/I}(z)=\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}.$
@@
It is clear that $(A)\Rightarrow(B)\Rightarrow(C)\Rightarrow(D)$. When $r\leq 1$, it follows from (2.1.1) that $(D)\Rightarrow(C)$ holds. Also when $r\neq 1$, we have $d_{\rm reg}(I)=\widetilde{d}_{\rm reg}(I)$, whence $(C)\Rightarrow(B)$ holds. For $m\leq n$, Proposition 2.2.4 implies $(B)\Rightarrow(C)$. ∎
2.3 Genericness and a variant of Fröberg’s conjecture
Given $m$ non-negative integers $d_{1},\ldots,d_{m}$, let $V(n,m,d_{1},\ldots,d_{m})\!=\!R_{d_{1}}\times\cdots\times R_{d_{m}}$, which is topologically identified with the affine space $\mathbb{A}^{N_{1}}(K)\times\dots\times\mathbb{A}^{N_{m}}(K)$ of dimension $\sum_{j=1}^{m}N_{j}$ with $N_{j}:=\binom{n+d_{j}-1}{d_{j}}$. Here the identification from $R_{d}$ to $\mathbb{A}^{N}$, where $N=\binom{n+d-1}{d}$, is given as follows: Letting $\{t_{1},\ldots,t_{N}\}$ be the set of all monomials in $R$ of degree $d$, each polynomial $f=\sum_{i=1}^{N}c_{i}t_{i}$ in $R_{d}$ is identified with its coefficients vector $(c_{1},\ldots,c_{N})$. We write simply $V$ for $V(n,m,d_{1},\ldots,d_{m})$ when we fix the parameter $(n,m,d_{1},\ldots,d_{m})$.
Definition 2.3.1 (cf. [34, Section 1]).
For given $n$, $m$, and $(d_{1},\ldots,d_{m})$, a property $\mathcal{P}$ for $\bm{F}=(f_{1},\ldots,f_{m})\in V$ is said to be generic (or, we say that a generic sequence in $V$ satisfies $\mathcal{P}$) if it holds on a non-empty Zariski open set in $V$, namely there exists a non-empty Zariski open set $U$ in $V$ such that
$U\subset\{\bm{F}=(f_{1},\ldots,f_{m})\in V:\mbox{$\bm{F}$ satisfies $\mathcal{P}$}\}.$

### v2

@@ 1 Introduction
  • The sequence $(f_{1}^{h},\ldots,f_{m}^{h})\in(R^{\prime})^{m}$ is generalized cryptographic semi-regular.
Then, it follows from Theorem 1.1 that $\max\{d_{\rm reg}(\langle F^{\rm top}\rangle),\widetilde{d}_{\rm reg}(\langle F^{h}\rangle)\}\leq D^{(n+1,m)}$ with $F^{\rm top}:=\{f_{1}^{\rm top},\ldots,f_{m}^{\rm top}\}$ and we have the complexity bound
$O\left(m\binom{n+D^{(n+1,m)}}{D^{(n+1,m)}}^{\omega}\right).$ (1.6)
Note that this upper-bound is heuristically found in [28].
@@ 2 Preliminaries
Let $K$ be a field, and let $\overline{K}$ denote its algebraic closure. Let $n$ be a natural number, and let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring of $n$ variables over $K$. We denote by $\mathcal{T}$ the set of monomials in $\{x_{1},\ldots,x_{n}\}$. As our convention, as graded structure objects, for a subset $S$ of $R$ and a non-negative integer $d$, we also denote by $S_{d}$ the subset consisting of all elements of degree $d$ in $S$. As an exception, if $S$ is an ideal of $R$, it is allowed that $S_{d}$ contains 00. The Krull dimension of a commutative ring $A$ is denoted by $\mathrm{Krull.dim}(A)$. For ideals $I$ and $J$ of $R$, their ideal quotient is defined as $(I:J)=\{f\in R\mid fJ\subset I\}$. The saturation of $I$ with respect to $J$ is $(I:J^{\infty}):=\{f\in R\mid fJ^{s}\subset I\mbox{ for some $s\geq 1$}\}$. In particular, the saturation of a homogeneous ideal $I$ with respect to the maximal homogeneous ideal $\mathfrak{m}=\langle x_{1},\ldots,x_{n}\rangle_{R}$ is denoted by $I^{\rm sat}=(I:\mathfrak{m}^{\infty})$, which is also homogeneous.
@@ 2.1 Hilbert regularity
Let $I$ be a non-zero homogeneous ideal of $R$. Let ${\rm HP}_{R/I}\in\mathbb{Q}[z]$ denote the Hilbert polynomial of $R/I$, namely there exists a non-negative integer $d_{0}$ such that ${\rm HF}_{R/I}(d)={\rm HP}_{R/I}(d)$ for any $d$ with $d\geq d_{0}$, where $\mathrm{HF}$ is the Hilbert function.
Definition 2.1.1 (Hilbert regularity (or index of regularity [14, Chapter IX, Section 3])).
@@
    There exists a non-negative integer $d_{0}$ such that ${\rm HF}_{R/I}(d)={\rm HF}_{R/I}(d_{0})>0$ (resp. ${\rm HF}_{R/I}(d)=0$) for any $d$ with $d\geq d_{0}$.
@@
The middle and the right hand sides are called the generalized degree of regularity of $I$ for $r\leq 1$, denoted by $\widetilde{d}_{\rm reg}(I)$ (cf. [36, Sect. 2.3]): When $r=0$, these are equal to the degree of regularity $d_{\rm reg}(I):=\min\{d:R_{d}=I_{d}\}$. Namely, we have
@@
Note also that the Hilbert series ${\rm HS}_{R/I}$ is written as ${\rm HS}_{R/I}(z)=h_{R/I}(z)/{(1-z)^{r}}$ for some polynomial $h_{R/I}(z)\in\mathbb{Z}[z]$, which is called the $h$-polynomial of $R/I$. Then, we have ${\rm hilb}(I)=\deg h_{R/I}-r+1$, see [7, Prop. 4.1.12].
@@
where ‘${\rm reg}$’ means Castelnuovo-Mumford regularity, see [17, §20.5] for the definition. The equality holds if $R/I$ is Artinian, i.e., the Krull dimension of $R/I$ is zero, see e.g., [18, Corollary 4.4].
Definition 2.1.4 (cf. [31, Section 3]).
@@
Lemma 2.1.5 ([6, Lemma 4.1]; see also [40, Proposition 5.4]).
Let $I$ be a homogeneous ideal of $R$, and let $\prec$ be the graded reverse lexicographical ordering on $R$ with $x_{n}\prec\cdots\prec x_{1}$. Then, the following are equivalent:
@@
    $\mathrm{Krull.dim}(R/(I+\langle x_{n-r+1},\ldots,x_{n}\rangle))=0$.
@@
    $\mathrm{Krull.dim}(R/(\langle\mathrm{LM}_{\prec}(I)\rangle+\langle x_{n-r+1},\ldots,x_{n}\rangle))=0$.
In the proof of [6, Lemma 4.1], the authors first proved the equivalence of the conditions (2), (3), and (4), and the implication (1) $\Rightarrow$ (2). The non-trivial part is that (2), (3), and (4) imply (1), which follows from [40, Proposition 5.4].
@@
${(I+\langle x_{n-i+1},\ldots,x_{n}\rangle)^{\rm sat}}=({(I+\langle x_{n-i+1},\ldots,x_{n}\rangle)}:x_{n-i}^{\infty})$
@@
For equivalent definitions, see [30, Section 2] for details. Note that $I$ is in SNP if and only if $\langle\mathrm{LM}_{\prec}(I)\rangle$ is in SNP (cf. [30, Proposition 2.11]). Hashemi also proved the following inequality:
Theorem 2.1.7 ([30, Proposition 4.7 and Theorem 4.17]).
@@
for the graded reverse lexicographical ordering $\prec$ on $R$ with $x_{n}\prec\cdots\prec x_{1}$.
Lemma 2.1.8 ([31, Lemma 23] and [32, Proposition 3.2 (v)]).
@@
    $I$ is in quasi stable position with respect to the ordered variables $x_{1},\ldots,x_{n}$ (see e.g., [31, Definition 6] for definition).
@@
The equivalence of (1) and (2) is proved in [31, Lemma 23], and that of (2) and (3) follows from [32, Proposition 3.2 (v)]. ∎
@@
Theorem 2.1.9 ([31, Proof of Theorem 30]).
Assume that $r=1$. If $R/(I+\langle x_{n}\rangle)$ is Artinian (or one of the equivalent conditions given in Lemma 2.1.8 holds), then we have the inequality (1.2) for the graded reverse lexicographical ordering $\prec$ on $R$ with $x_{n}\prec\cdots\prec x_{1}$.
Here, we recall a fact that, if $K$ contains sufficiently many elements, the assumption $\mathrm{Krull.dim}(R/(I+\langle x_{n}\rangle))=0$ in Theorem 2.1.9 is satisfied after a suitable linear changer of coordinates. In the rest of this paper, we explain this fact in details, since we could not find any appropriate reference.
Although it may be an easy consequence of the theory of hyperplane sections, there exists a hyperplane defined over $K$ that does not pass through any of given finite points in a projective space over $\overline{K}$ if $K$ is large enough. Here, we prove a somehow detailed argument:
Lemma 2.1.10.
Let $\bm{p}_{1},\ldots,\bm{p}_{N}$ be mutually distinct points in $\mathbb{P}^{n-1}(\overline{K})$. If the cardinality of $K$ is greater than $N$, then there exists a linear form $\ell\in R$ that does not vanish at any of $\bm{p}_{1},\ldots,\bm{p}_{N}$.
Proof.
For each point $\bm{p}_{k}=(p_{1,k}:\cdots:p_{n,k})\in\mathbb{P}^{n-1}(\overline{K})$, we denote by $S_{k}$ the set of all linear forms that vanish at $\bm{p}_{k}$, say
$S_{k}=\{y_{1}x_{1}+\cdots+y_{n}x_{n}\mid y_{i}\in K\mbox{ and }y_{1}p_{1,k}+\cdots{+}y_{n}p_{n,k}=0\},$
$S_{k}^{\prime}=\{(y_{1},\ldots,y_{n})\in K^{n}\mid{p_{1,k}y_{1}+\cdots{+}p_{n,k}y_{n}=0\}}.$
The dimension of $S_{k}^{\prime}$ as a $K$-linear space is at most $n-1$. Indeed, for an arbitrary non-zero entry $p_{i,k}$ of $\bm{p}_{k}$, the vector with $1$ in the $i$-th entry and 00’s elsewhere does not belong to $S_{k}^{\prime}$, so that $S_{k}^{\prime}$ is a proper subspace of $K^{n}$.
When $K$ is an infinite field, it follows from the irreducibility of an affine space over $K$ that $\bigcup_{k=1}^{N}S^{\prime}_{k}$ cannot cover $K^{n}$. When $K$ is a finite field of order $q$, we have
#${\color[rgb]{0,0,0}\#\left(\bigcup_{k=1}^{N}S_{k}^{\prime}\right)\leq\sum_{k=1}^{N}\#S_{k}^{\prime}\leq N\cdot q^{n-1}<q^{n}},$
whence a desired linear form over $K$ exists. ∎
Lemma 2.1.11.
With notation as above, assume that $\mathrm{Krull.dim}(R/I)\leq 1$. If the cardinality of $K$ is greater than #$\#V_{\overline{K}}(I)$, then there exists a linear form $\ell\in R$ such that $\mathrm{Krull.dim}(R/(I+\langle\ell\rangle))=0$. Moreover, there also exists a suitable linear coordinate change $\sigma$ sending $\ell$ to $x_{n}$ such that $\mathrm{Krull.dim}(R/(I^{\sigma}+\langle x_{n}\rangle))=0$.
Proof.
The first part follows from Lemma 2.1.10. There exists a $k\in\{1,\ldots,n\}$ such that the $x_{k}$-coefficient of $\ell$ is not zero. We may also assume that the $x_{k}$-coefficient is $1$, without loss of generality. Let $\sigma_{1}$ be the linear transformation of variables $x_{1},\ldots,x_{n}$ defined by a permutation exchanging $x_{k}$ and $x_{n}$, and write the image of $\ell$ by $\sigma_{1}$ as $\ell^{\sigma_{1}}=a_{0}x_{0}+a_{1}x_{1}+\cdots+a_{n}x_{n}$ for $a_{i}\in K$ with $a_{n}=1$. Then, let $\sigma_{2}$ be a linear transformation of variables $x_{1},\ldots,x_{n}$ defined by
$h^{\sigma_{2}}:=h(x_{1},\ldots,x_{n-1},x_{n}-a_{1}x_{1}-\cdots-a_{n-1}x_{n-1})$
for $h\in R$. Then, the invertible transformation $\sigma:=\sigma_{2}\circ\sigma_{1}$ of variables $x_{1},\ldots,x_{n}$ sends $\ell$ to $x_{n}$, so that $\mathrm{Krull.dim}(R/(I^{\sigma}+\langle x_{n}\rangle))=0$. ∎
Remark 2.1.12.
In Lemma 2.1.11, we have an exact sequence
${(R/I)_{d^{\prime}-1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\times\ell}$${(R/I)_{d^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(R/(I+\langle\ell\rangle))_{d^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
for each $d^{\prime}\geq 1$. Therefore, the minimal $d$ for which the $-\ell$ map $(R/I)_{d^{\prime}-1}\to(R/I)_{d^{\prime}}$ is surjective for any $d^{\prime}\geq d$ is equal to $d_{\rm reg}(I+\langle\ell\rangle)$.
@@ 2.2 Generalized semi-regular sequences
Definition 2.2.2 ([2, Definition 3]; see also [15, Definition 1]).
@@
Definition 2.2.3 ([2, Definition 5], [4, Definition 5]; see also [15, Section 2]).
@@
Proposition 2.2.4 ([15, Proposition 1 (d)]; see also [4, Proposition 6]).
With the same notation as in Theorem 2.2.1, a sequence $(f_{1},\ldots,f_{m})$ of homogeneous polynomials in $R\smallsetminus K$ is cryptographic semi-regular if and only if
@@
Remark 2.2.5.
In [25, Definition 1.3], Fröberg-Hollman already defined a sufficiently generic ideal as one generated by homogeneous polynomials $f_{1},\ldots,f_{m}\in R$ satisfying (2.2.2). Namely, a cryptographic semi-regular sequence is nothing but a sequence of homogeneous polynomials generating a sufficiently generic ideal.
Here, we recall the following lemma by Fröberg:
Lemma 2.2.6 ([23, b) on page 121]).
@@
in the lexicographic sense, i.e., the first coefficient that differs, if any, is greater on the left hand side, see [23, page 118].
@@
To deal with the case where $R/I$ has Krull dimension one, we extend the notion of cryptographic semi-regular as follows:
Definition 2.2.8 ([36, Definition 2.3.3]).
@@
Remark 2.2.9 (cf. [36, Section 4.1]).
A generalized cryptographic semi-regular (but neither regular nor semi-regular) sequence is easily generated. For example, choosing coefficients of $g_{1},\ldots,g_{m}$ with $m\geq n$ uniformly at random from a finite field $K=\mathbb{F}_{q}$ and choosing an arbitrary point $\bm{p}=(p_{1},\ldots,p_{n-1},p_{n})\in\mathbb{A}^{n}(K)$ with $p_{n}=1$, then it is expected that $(f_{1},\ldots,f_{m})$ with $f_{j}:=g_{j}(x_{1},\ldots,x_{n})-g_{j}(p_{1},\ldots,p_{n})x_{n}^{d_{j}}$ is generalized cryptographic semi-regular (but neither regular nor semi-regular). Without loss of generality, we may take $\bm{o}=(0:\cdots:0:1)$ as $\bm{p}$, see Section 4 below for details.
Example 2.2.10.
Let $K=\mathbb{F}_{31}$, $n=3$, and $R=K[x_{1},x_{2},x_{3}]$. Here we consider the case where $(d_{1},d_{2},d_{3})=(3,2,2)$ with $m=3$, and let
$f_{1}$ $=$ $27x_{1}^{3}+5x_{1}^{2}x_{2}+6x_{1}x_{2}^{2}+20x_{2}^{3}+29x_{1}x_{2}x_{3}+25x_{2}^{2}x_{3}+3x_{1}x_{3}^{2}+26x_{2}x_{3}^{2},$
$f_{2}$ $=$ $2x_{1}^{3}+29x_{1}^{2}x_{2}+25x_{1}x_{2}^{2}+3x_{2}^{3}+25x_{1}^{2}x_{3}+20x_{1}x_{2}x_{3}+25x_{2}^{2}x_{3}+19x_{1}x_{3}^{2}+24x_{3}x_{3}^{2},$
$f_{3}$ $=$ $12x_{1}^{2}+12x_{1}x_{2}+23x_{2}^{2}+x_{1}x_{3}+10x_{2}x_{3}.$
Putting $I=\langle f_{1},f_{2},f_{3}\rangle$, one can verify that
${\rm HS}_{R/I}(z)=1+3z+5z^{2}+5z^{3}+3z^{4}+z^{5}+z^{6}+\cdots$
with $\mathrm{Krull.dim}(R/I)=1$, $D:=\mathrm{hilb}(I)=5$, and $\dim_{K}(R/I)_{D}=1$, so that $(f_{1},f_{2},f_{3})$ is $D$-regular (in fact $D+1$-regular). Namely $(f_{1},f_{2},f_{3})$ is generalized cryptographic semi-regular, but is neither regular nor cryptographic semi-regular.
@@
By our assumption that $\bm{F}$ is generalized cryptographic semi-regular, the equality (2.2.1) holds for $d=\widetilde{d}_{\rm reg}(I)$. For any $d^{\prime}$ with $d^{\prime}<\widetilde{d}_{\rm reg}(I)$, we have ${\rm HF}_{R/I}(d^{\prime})>0$, whence the $z^{d^{\prime}}$-coefficient of $(\prod_{j=1}^{m}(1-z^{d_{j}}))/(1-z)^{n}$ is also positive. Thus, the inequality (2.2.3) holds. ∎
@@
$\bm{F}$ is regular, equivalently ${\rm HS}_{R/I}(z)=(\prod_{j=1}^{m}(1-z^{d_{j}}))/(1-z)^{n}$.
@@
It is clear that $(A)\Rightarrow(B)\Rightarrow(C)\Rightarrow(D)$. When $r\leq 1$, it follows from (2.1.1) that $(D)\Rightarrow(C)$ holds. Also when $r\neq 1$, we have $d_{\rm reg}(I)=\widetilde{d}_{\rm reg}(I)$, whence $(C)\Rightarrow(B)$ holds. For $m\leq n$, Proposition 2.2.4 implies $(B)\Rightarrow(C)$.
On the other hand, it is clear that (B) does not imply (A) if $m>n$. The sequence $(f_{1},f_{2},f_{3})$ provided in Example 2.2.10 gives a counterexample for the implication (C) $\Rightarrow$ (B) when $r=1$. As for (D) $\Rightarrow$ (C), here is a counterexample with $r=2$: Let $(g_{1},g_{2}):=(x^{3},xy)\in\mathbb{Q}[x,y,z]^{2}$ and put $I:=\langle g_{1},g_{2}\rangle$. Then one can verify that $\mathrm{HS}_{R/I}(z)=1+3z+5z^{2}+6z^{3}+7z^{4}+\cdots$ with $\mathrm{Krull.dim}(R/I)=2$ and $D:=\mathrm{hilb}(I)=2$, whence $(g_{1},g_{2})$ is $D$-regular but not regular. ∎
2.3 Genericness and Fröberg’s conjecture
Given $m$ positive integers $d_{1},\ldots,d_{m}$, let ${V_{n,m,(d_{1},\ldots,d_{m})}}:=R_{d_{1}}\times\cdots\times R_{d_{m}}$, which is topologically identified with the affine space $\mathbb{A}^{N_{1}}(K)\times\dots\times\mathbb{A}^{N_{m}}(K)=\mathbb{A}^{N}(K)$ of dimension $N:=\sum_{j=1}^{m}N_{j}$ with $N_{j}:=\binom{n+d_{j}-1}{d_{j}}$. Here, we identify a form in $R_{d_{j}}$ with a point in $\mathbb{A}^{N_{j}}(K)$ by taking the coefficients. We may write $K^{N}$ simply for $\mathbb{A}^{N}(K)$. We also correspond a point $\bm{F}=(f_{1},\ldots,f_{m})\in V_{n,m,(d_{1},\ldots,d_{m})}$ to an ideal $\langle\bm{F}\rangle_{R}:=\langle f_{1},\ldots,f_{m}\rangle_{R}$, where the subscript $R$ may be omitted if the meaning is clear from the context.
Following Fröberg’s definition (cf. [23], [24], [27]), we say that a form of $R_{d}$ is generic if all monomials of degree $d$ occur with nonzero coefficient and if all the coefficients are algebraically independent over the prime field of $K$. We call $\bm{F}=(f_{1},\ldots,f_{m})\in V_{n,m,(d_{1},\ldots,d_{m})}$ a generic sequence if all $f_{i}$’s are generic forms with coefficients that are mutually algebraic independent over the prime field. If $K$ does not contain $N$ elements that are algebraically independent over the prime field of $K$ (for example $K$ is an algebraic extension of the prime field), we may replace $K$ by its arbitrary extension field $L$ containing such $N$ elements. When we consider a property of Hilbert series, this makes a sense since every Hilbert series is not changed by extending the field of coefficients.
On the other hand, the genericness of a property for $\bm{F}\in V_{n,m,(d_{1},\ldots,d_{m})}$ is defined as follows:
Definition 2.3.1 (cf. [43, Section 1]).
For given $n$, $m$, and $(d_{1},\ldots,d_{m})$, a property $\mathcal{P}$ for $\bm{F}=(f_{1},\ldots,f_{m})\in V_{n,m,(d_{1},\ldots,d_{m})}$ is said to be generic if it holds on a non-empty Zariski open set in $V_{n,m,(d_{1},\ldots,d_{m})}$, namely there exists a non-empty Zariski open set $U$ in $V_{n,m,(d_{1},\ldots,d_{m})}$ such that
$U\subset\{\bm{F}=(f_{1},\ldots,f_{m})\in V_{n,m,(d_{1},\ldots,d_{m})}:\mbox{$\bm{F}$ satisfies $\mathcal{P}$}\}.$
Here, we recall Fröberg’s important lemma for the minimality of the Hilbert series on generic sequences. Note that the characteristic of $K$ is assumed to be 00 in the original paper [23], but it can be easily proved that the assertion of the lemma holds for any characteristic. We here state the lemma in a slightly modified (but equivalent) form for the convenience of our use.
Lemma 2.3.2 ([23, Lemma 1]).
With notation as above, let $\Omega$ be an extension field of $K$, and let $\mathcal{C}:=\{c_{j,t}:1\leq j\leq m,\ t\in\mathcal{T}_{d_{j}}\}$ be a set of elements in $\Omega$ that are algebraically independent over the prime field $K_{0}$ of $K$. Here $\mathcal{T}_{d}$ is the set of monomials of degree $d$ in $\{x_{1},\ldots,x_{n}\}$. Let $L:=K(\mathcal{C})$, let $R_{L}:=R\otimes_{K}L=L[x_{1},\ldots,x_{n}]$, and put $g_{j}:=\sum_{t\in\mathcal{T}_{d_{j}}}c_{j,t}t\in(R_{L})_{d_{j}}$ for each $1\leq j\leq m$. Namely $\bm{G}=(g_{1},\ldots,g_{m})$ is a generic sequence. Then, for any $\bm{F}=(f_{1},\ldots,f_{m})\in V_{n,m,(d_{1},\ldots,d_{m})}$, we have the coefficient-wise inequality
$\mathrm{HS}_{R/\langle f_{1},\ldots,f_{m}\rangle_{R}}(z)=\mathrm{HS}_{R_{L}/\langle f_{1},\ldots,f_{m}\rangle_{R_{L}}}\geq\mathrm{HS}_{R_{L}/\langle g_{1},\ldots,g_{m}\rangle_{R_{L}}}(z).$ (2.3.1)